OpenStudy (anonymous):

I need help in solving. 3/5x - 1/16 = 4/9

OpenStudy (anonymous):

is that 3/5 times x , or 3 divided by 5x

OpenStudy (anonymous):

times

OpenStudy (anonymous):

it is a fraction

OpenStudy (anonymous):

ok if its 3/5 times x then one way to proceed is to multiply each term by lcd of 5,16 and 9

OpenStudy (anonymous):

sorry must go

jimthompson5910 (jim_thompson5910):

\[\large \frac{3}{5}x-\frac{1}{16}=\frac{4}{9}\] \[\large 720\left(\frac{3}{5}x-\frac{1}{16}\right)=720\left(\frac{4}{9}\right)\] \[\large 432x-45=320\] \[\large 432x=320+45\] \[\large 432x=365\] \[\large x=\frac{365}{432}\]

jimthompson5910 (jim_thompson5910):

In the second step, I multiplied both sides by the LCD 720 to clear out the fractions.

OpenStudy (anonymous):

ty